Give a contrapositive proof of the theorem; "If n is an interfer and 3n + 2 is even, then n is even."?


1
Expert's answer
2021-05-06T15:02:30-0400

Let us give a contrapositive proof of the theorem: "If "n" is an interger and "3n + 2" is even, then "n" is even.".

For this suppose that "n" is an interger and "3n + 2" is even, but "n" is not even, and hence "n" is odd. Then "n=2k+1" for some "k\in\mathbb Z." It follows that "3n+2=3(2k+1)+2=6k+3+2=(6k+4)+1=2(3k+2)+1", where "3k+2" is integer. Therefore, "3n+2" is odd. We conclude that "3n+2" is not even, that is we have a contradiction with initial assumption that "3n + 2" is even.